OpenStudy (anonymous):

simplify (sinx)/(1+cosx)-(1+cosx)/(sinx)

OpenStudy (queelius):

\[\frac{ \sin(x) }{ 1+\cos(x) } - \frac{ 1 + \cos(x) }{ \sin(x) }\] So, let's add the terms together. To do this, we need to make them have the same denominator. \[\frac{ \sin(x)\sin(x) }{ (1+\cos(x))(\sin(x)) }-\frac{ (1+\cos(x))(1+\cos(x)) }{ \sin(x)(1+\cos(x)) }\] Adding them together we get. \[\frac{ \sin(x)\sin(x)-(1+\cos(x))(1+\cos(x)) }{ (1+\cos(x))\sin(x) }\] Now, there are some simplifications we can do. \[\frac{ \sin(x)\sin(x)-1-2\cos(x)+\cos(x)\cos(x) } { (1+\cos(x))\sin(x) }\] From the unit circle, we also know that sin(x) * sin(x) + cos(x) * cos(x) = 1. So, what's next?

OpenStudy (anonymous):

Ty.

OpenStudy (anonymous):

I have more problems, but that was the one that I did not understand on that assignment.

OpenStudy (queelius):

It's not done yet. :)

OpenStudy (anonymous):

-2cosx/(1+cosx)sinx

OpenStudy (queelius):

Yup. There are other things you can do to it, but I don't necessarily think they're simplifications. Maybe you can spot some simplifications I can't thought.

OpenStudy (anonymous):

Thanks, I'll send you a message if I get another question.

OpenStudy (queelius):

Ok, glad I could help.